OpenStudy (anonymous):

how would you factor this: y(squared)+8y-20

OpenStudy (anonymous):

y^2+8y-20 (y+10)(y-2)

OpenStudy (anonymous):

y^2 + 8y -20 = y^2 + 10y -2y -20 = y(y+10) -2(y+10) = (y+10) (y-2)